OpenStudy (anonymous):

solve the equation 3x+6=x-8

OpenStudy (poofypenguin):

3x+6=x-8 3x-x=-8-6 2x=-14 x=(-14/2) x=-7

OpenStudy (anonymous):

ooh I just love doin this kind of question but @PoofyPenguin got the answer already XD

OpenStudy (poofypenguin):

I love these kind of questions too! I guess we have that in common, lol!! sorry @Trexy!!! :D

OpenStudy (anonymous):

its fine.. hahah.. were just helping people XD

OpenStudy (anonymous):

@Crash2 hope you understand the problems